Determine the area under the standard normal distribution curve between -0.91 and 2.5.


1
Expert's answer
2022-04-26T11:55:33-0400

"P(-0.91<Z<2.5)=P(Z<2.5)-P(Z<-0.91)="

"=P(Z<2.5)-1+P(Z<0.91)=0.9938-1+0.8186=0.8124."